OpenStudy (anonymous):

Let f(x)=x^2-3 and h(x)=x+3. Find the following value or expression (foh)(5).

OpenStudy (anonymous):

@FibonacciChick666

OpenStudy (kc_kennylau):

\[\Large f(x)=x^2-3\\\Large h(x)=x+3\\\Large (f\circ h)(5) = f(h(5))\]

OpenStudy (kc_kennylau):

Let's plug \(\Large5\) into \(\Large h\)

OpenStudy (anonymous):

ok so \[(foh)(5)=f(h(5))\] \[f(5+3)=f(8)\]

OpenStudy (kc_kennylau):

Yes

OpenStudy (anonymous):

then \[f(8)= (8)^2-3\]

OpenStudy (anonymous):

equals 61

OpenStudy (unklerhaukus):

good work!

OpenStudy (anonymous):

I'm not sure why I needed help with that one....Maybe I just needed someone to watch me LOL! Happens...thanks guys :)